OpenStudy (anonymous):

Please help me with this one, I'm struggling with it. Integral of 1 / x^2.(x+1)

OpenStudy (amistre64):

partial decomp?

OpenStudy (amistre64):

it splits into linear denominators ... \[\frac{1}{x^2(x+1)}=\frac{A}{x}+\frac{B}{x^2}+\frac{C}{x+1}\]

OpenStudy (anonymous):

What if I put like this = A/(x+1) + (Bx+C)/x^2

OpenStudy (amistre64):

then its wrong ...

OpenStudy (amistre64):

you have to consider all the possible ways that the decomp can occur

OpenStudy (anonymous):

lol I guess. Thanks a lot

OpenStudy (anonymous):

After calculating, i got A=0, B=1, C=0. does it sound right?

OpenStudy (amistre64):

the x^2 denom always makes me uneasy :) it looks quadratic to me and I always want to put a linear form above it ... but for some reason its assume to be a multiplicity of a single linear and a constant above it is stated in the texts

OpenStudy (amistre64):

i dont get those values ... can you show your work?

OpenStudy (anonymous):

ok. When I redo it, I got: 1 = Ax(x+1) + B(x+1) + Cx^2 1 = Ax^2 + Ax + Bx + B + Cx^2 1 = x^2(A+C) + x(A+B) + B => B=1 A + B= 0 => A=-1 A + C = 0 => C = 1

OpenStudy (amistre64):

there are 2 methods that I recall, a coverup and a rework. the coverup is fine for some cases but not all cases so I just go with the rework .. try to rebuild how you would add them: \[\frac{1}{x^2(x+1)}=\frac{A}{x}+\frac{B}{x^2}+\frac{C}{x+1}\] \[\frac{1}{x^2(x+1)}=\frac{Ax(x+1)}{x~x(x+1)}+\frac{B(x+1)}{x^2(x+1)}+\frac{Cx^2}{x^2(x+1)}\] \[\frac{1}{x^2(x+1)}=\frac{Ax(x+1)+B(x+1)+Cx^2}{x^2(x+1)}\] these are only equal if the parts are equal to one another giving us the tops to compare \[1=Ax(x+1)+B(x+1)+Cx^2\] lets zero out the terms when x=0 we are left with:\[1=B(0+1)\] when x=-1 we are left with:\[1=C(-1)^2\] knowing B and C we can assume any value for x to assess the A, lets use something simple like x=1: \[1=A(1)(1+1)+1(1+1)+1(1)^2\]

OpenStudy (amistre64):

A = -3/2 ... if im seeing it right

OpenStudy (anonymous):

@amistre64, the decompositions are the same \[\frac{A}{x+1}+\frac{Bx+C}{x^2}=\frac{A}{x+1}+\frac{B}{x}+\frac{C}{x^2}\]

OpenStudy (amistre64):

yeah, hartnn pointed that out to me in a different question :) i either had a bad textbook, or a truamatic memory lol